OpenStudy (anonymous):

A=4200[1+ 7/100]^(12)

OpenStudy (anonymous):

i cant seem to figure this out, can someone help me please?

OpenStudy (anonymous):

this is my last question and im taking a long time on it i want to go to sleep lol

OpenStudy (anonymous):

just do it??

OpenStudy (anonymous):

i dont know how i am math illiterate

OpenStudy (anonymous):

so serious....

OpenStudy (anonymous):

take 1+7/100 to the 12 power

OpenStudy (anonymous):

then multiplu by 4200

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A=4200(107/100)^12 that would be A=4200(1.07)^12 that is 9459.2046 ....... ok ( from calculator)

OpenStudy (anonymous):

ok thx i really appreciate it .

OpenStudy (anonymous):

yeah i dont have my calc on me

OpenStudy (anonymous):

thx everyone for helping me i really appreciate it :)))